directions: set up and solve the following.
question: a change purse contains 14 coins consisting of nickels and dimes and having a total value of $1.10. how many nickels are in the change purse?

Found 2 solutions by stanbon, andkon:
Answer by stanbon(48522)

question: a change purse contains 14 coins consisting of nickels and dimes and having a total value of $1.10. how many nickels are in the change purse?
---------------
Let the number of nickels be "x"; value of these is 5x cents
the number of dimes is "14-x" ; value of these is 10(14-x)=140-10x cents
---------------------------
EQUATION:
value + value = 110 cents
5x + 140-10x = 110
-5x = -30
x= 6 (number of nickels)
14-x= 8 (number of dimes)
===================
Cheers,
Stan H.

Answer by andkon(15)